OpenStudy (anonymous):

3. A rectangular classroom is twice as long as it is wide, and its area is 800 square feet. Find the dimensions of the classroom.

OpenStudy (anonymous):

800 = 2x(x) 800 = 2x^2 0 = 2x^2 - 800 0 = 2(x^2 - 400) Can you solve the quadratic now?

OpenStudy (anonymous):

whould it be 2(x+2)(x-200)

OpenStudy (anonymous):

No. Do you recognize that it's a difference in squares pattern? \[a^{2} - b^{2} = (a - b)(a + b)\]

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

now i do

OpenStudy (anonymous):

would the answer be 2(x-20)(x+20) X=20 X=-20

OpenStudy (anonymous):

There you go.

OpenStudy (anonymous):

Do yuo understand that dimensions can't be negative? Therefore, the width is 20. Then, the length is twice as long as the width, so 2(20) = 40 40 is the length. 20 is the width.

OpenStudy (anonymous):

yes thank you, do you think you could help me with a few others

OpenStudy (anonymous):

A plane is flying nonstop from Pittsburgh to Los Angles, a distance of about 2400 miles. After 1 hour 15 minutes in the air, the plane flies over St Louis (a distance of 800 miles from Pittsburgh). Estimate the time the plane will arrive in Los Angles if it departs from Pittsburgh at 12 p.m.(noon).